# DeMPT: Decoding-enhanced Multi-phase Prompt Tuning for Making LLMs Be Better Context-aware Translators

### v1

@@ 3.2 Experimental Results
The main experimental results are presented in Tables 1 and 2. Additionally, a comparison of the number of trainable parameters is presented in Table 3 across different tuning methods. From these results we have the following observations:
The encoder-decoder Transformer model (Trans.) performs better than the LLMs with LoRA tuning in most translation directions in BLEU score. For example, when utilizing llama-2-7b as the foundation model, Transformer surpasses MT-LoRA an average of 0.75 BLEU score across all translation tasks. However, MT-LoRA model outperforms Trans. in terms of COMET, suggesting that translations from LLMs may align more closely with human preferences.
The MT-PT model presents superior performance compared to the MT-LoRA model in terms of BLEU, COMET, and BlonDe. Taking bloomz-7b1-mt as the foundation model, the MT-PT model outperforms the MT-LoRA model by an average of 3.84 BLEU score, 3.51 BlonDe score, and 0.0047 COMET score. Nevertheless, the MT-PT model sacrifices efficiency for performance, introducing more trainable parameters (13.87% vs. 0.12%).
Leveraging the inter-sentence context is helpful in alleviating discourse-related issues. For example, with bloomz-7b1-mt used as the foundation model, the CMT-PT model, despite underperforming in BLEU and COMET compared to the MT-PT model, excels in discourse-related BlonDe scores (averaging 57.66 vs. 57.17).
Our MPT/DeMPT model outperforms all baselines across all translation tasks. For example, when using llama-2-7b as the foundation model, our MPT model achieves an average gain of 1.62/1.45/2.03 in BLEU/COMET/BlonDe score compared to the CMT-PT model. Furthermore, our decoding-enhance strategy enhances the capacity of LLMs in context-aware NMT, with DeMPT outperforming MPT in BLEU/COMET/BlonDe score (averaging 42.39/0.8758/59.68 vs. 41.74/0.8716/58.91).
The model built upon llama-2-7b as the foundation model outperforms the one using bloomz-7b1-mt, suggesting that llama-2-7b serves as a more robust foundation model for translation tasks.
| MT-LoRA | MT-PT/CMT-PT | DeMPT | |
@@ 4 Discussion
In this section, we use bloomz-7b1-mt as the foundation model to discuss and analyze our approach. See Appendix CE for further discussions.
@@ 4.1 Effect of Length of Inter-sentence Context
For efficient training, we define the inter-sentence context in Section 2 as previous sentences with a total tokens not exceeding 256. We are curious about the potential impact of inter-sentence length on the performance of our approach. Consequently, we extend the inter-sentence context length from 256 to 1024 and assess the performance of our approach in the ZHEN task.
@@
Figure 4 shows the performance trend of the CMT-PT model and our DeMPT model. As the length of the inter-sentence context increases, both models exhibit a slight enhancement in both BLEU and BlonDe scores. Interestingly, our model with a 256-token inter-sentence context outperforms the CMT-PT model with a 1024-token inter-sentence context in both BLEU and BlonDe scores. This further suggests the effectiveness of our approach in harnessing the capabilities of LLMs for context-aware NMT compared to the concatenation strategy.
4.2 Effect of Prompt Length
As our approach is implemented based on deep prompt tuning, next we compare the impact of the trainable prompt length for MT-PT, CMT-PT, and our DeMPT.
Figure 5 shows the performance curves when increasing the prompt length from 32 to 128. We observe that increased prompt length tends to enhance performance for both BLEU and BlonDe, yet the gains exhibit diminishing returns. This finding is consistent with that in Li and Liang (2021); Lester et al. (2021); Tan et al. (2022). We also observe that DeMPT with a prompt length of 64 outperforms both MT-PT and CMT-PT with a prompt length of 128 on both metrics, suggesting the superiority of our approach over the concatenation strategy in enhancing LLMs’ capacity for context-aware NMT.
4.3 Comparison of Inference Speed
@@
4.4 Performance on Contrastive Test Set
We evaluate the models’ ability to resolve discourse inconsistencies using the contrastive test set proposed by Voita et al. (2019a), which focuses on four discourse phenomena such as deixis, lexicon consistency (lex.c), ellipsis inflection (ell.infl), and verb phrase ellipsis (ell.VP) in EnglishRussian translation. Within the test set, each instance comprises a positive translation and several negative ones that vary by only one specific word. The purpose of the contrastive test set is to assess whether a model is more inclined to generate a correct translation as opposed to incorrect variations.
Table 5 lists the accuracy of translation prediction on the contrastive test set for MT-PT, CMT-PT and DeMPT. Compared to the context-agnostic MT-PT model, both context-aware CMT-PT and DeMPT models show substantial improvements across the four discourse phenomena. Additionally, DeMPT demonstrates the best performance, surpassing CMT-PT by an average accuracy margin of 3.8.
@@ 4.4 Performance on Contrastive Test Set
4.5 Human Evaluation
We use the Direct Assessment (DA) method Graham et al. (2017) to manually assess the quality of translations generated by DeMPT and CMT-PT. In this assessment, human evaluators compare the meaning of the MT output with a human-produced reference translation, working within the same language.
Specifically, we randomly select 5 documents with a total of 200 groups of sentences from the ZHEN test set. To avoid potential bias in evaluation, we recruit 6 professional translators and ensure each translation from DeMPT or CMT-PT is scored twice by two translators. Table 6 shows the DA scores for CMT-PT and DeMPT. Our DeMPT outperforms CMT-PT by 7.14 DA score, providing strong evidence for the effectiveness of our approach. Further details and results regarding the DA can be found in Appendix C.
| Model | Score_1 | Score_2 | Average |
| CMT-PT | 79.00 | 80.17 | 79.59 |
| DeMPT | 86.17 (+7.17) | 87.30 (+7.13) | 86.73 (+7.14) |
5 Related Work
Large Language Models for Context-aware Machine Translation.
While traditional context-aware neural machine translation (NMT) has seen considerable progress in recent years Jean et al. (2017); Wang et al. (2017); Voita et al. (2018); Maruf et al. (2019); Kang et al. (2020); Bao et al. (2021); Sun et al. (2022); Bao et al. (2023), the effective integration of large language models (LLMs) to model inter-sentence context and enhance context-aware translation remains an area of limited exploration. Existing studies mainly focus on the assessment of LLMs’ ability in discourse modeling. For example, Wang et al. (2023) approach context-aware NMT as a task involving long sequence generation, employing a concatenation strategy, and conduct comprehensive evaluations of LLMs such as ChatGPT and GPT-4. Their focus includes the impact of context-aware prompts, comparisons with translation models, and an in-depth analysis of discourse modeling ability. Similarly, Karpinska and Iyyer (2023) engage professional translators to evaluate LLMs’ capacity in context-aware NMT. In contrast, Wu et al. (2024) compare the effectiveness of various parameter-efficient fine-tuning methods on moderately-sized LLMs for context-aware NMT. Besides, Wu and Hu (2023) explore the prompt engineering with GPT language models specifically for document-level (context-aware) MT while Li et al. (2024) experiment with combining sentence-level and document-level translation instructions of varying lengths to fine-tune LLMs.
Prompt Tuning for Large Language Model.
Liu et al. (2021) and Li and Liang (2021) propose to make LLMs adapt to various tasks by adding trainable prompts (also called continuous prompts) to the original input sequences. In this paradigm, only the continuous prompts are updated during training. Liu et al. (2022) further introduce deep prompt tuning, extending the idea by inserting trainable prompts into all layers of LLMs, rather than just the embedding layer. While these approaches lay the groundwork for a general framework, our focus lies in augmenting the performance of LLMs specifically for inter-sentence context modeling in context-aware NMT. Notably related, Tan et al. (2022) propose a multi-phase tuning approach to enhance the sentence-level translation performance of a multilingual GPT. However, the exploration of effective LLM tuning for addressing discourse-related challenges in the context-aware NMT domain remains underdeveloped.
6 Conclusion
Limitations
Owing to resource limitations, our work is restricted to moderate-scale LLMs, specifically those with 7 billion parameters, and a confined window size of inter-sentence context. It is imperative to acknowledge that the results of our research may differ when employing larger models and an extended window size for inter-sentence context. We acknowledge these limitations and consider them as avenues for future exploration.
Appendix A Datasets
Statistics and Splitting of Datasets.
We provide the detailed statistic in Table 7. For all translation tasks, we randomly select 80% document pairs from the corpus as the training set. Both the test set and validation set include 150 document pairs each, randomly sampled from the remaining 20% of document pairs in the corpus. Regarding sentence preprocessing across all datasets, we segment the sentences with the tokenizer from the respective foundation model. No additional preprocessing steps are performed. Datasets are downloaded from https://data.statmt.org/news-commentary/v18.
| Dataset | ZHEN | FREN | DEEN | ESEN | RUEN | |||||
| #Doc | #Sent | #Doc | #Sent | #Doc | #Sent | #Doc | #Sent | #Doc | #Sent | |
| Test | 150 | 5,747 | 150 | 5,795 | 150 | 5,967 | 150 | 5,819 | 150 | 5,619 |
Appendix B Training Details
For all Transformer NMT models, we use the transformer-base setting as in Vaswani et al. (2017), where the learning rate is set to 1e-4. The Transformer NMT models are trained on 4 NVIDIA V100 32GB GPUs with a batch size of 4096. For the models with prompt tuning in Section 3, including MT-PT, CMT-PT, and our MPT and DeMPT models, the length of the trainable prompt is set as 64. During both training and inference, the model generates only the current target sentence, operating in a many-to-one translation mode. For all fine-tuning models in this paper, we set the training epoch to 4, and the warm-up rate to 0.1. We use the log learning rate decay strategy with a maximum learning rate of 5e-5. We collate a mini-batch by counting the total tokens inside the batch and set the batch size as 4096. All fine-tuning models are trained on 4 NVIDIA A800 GPUs with Deespeed Zero 2 offload setting Rajbhandari et al. (2020).999https://github.com/microsoft/DeepSpeed
Appendix C Details of Human Evaluation

### v2

@@ 3.2 Experimental Results
The main experimental results are presented in Tables 1 and 2. Additionally, a comparison of the number of trainable parameters is presented in Table 3 across different tuning methods. When examining llama-2-7b and focusing on context-agnostic models, we find that the Transformer models (Trans.) generally outperform LLMs with LoRA tuning (MT-LoRA) in most translation directions based on BLEU score. However, the MT-LoRA models surpass Trans. in COMET, indicating that translations from LLMs may better align with human preferences. Additionally, the MT-PT models exhibit superior performance compared to the MT-LoRA models across BLEU, COMET, and BlonDe metrics. This improvement could be attributed to the more trainable parameters in the MT-PT models (13.87% vs. 0.12%).
Importantly, by comparing MT-PT and CMT-PT, we observe that CMT-PT which indiscriminately leverages the inter- and intra-sentence context with the concatenation way, even hurts performance for certain translation tasks. For example, the CMT-PT models, despite excelling in discourse-related BlonDe scores (averaging 57.65 vs. 57.17), underperforms in BLEU and COMET compared to the MT-PT models. In contrast, our context-aware MPT and DeMPT models outperform all LLM baselines across all translation tasks in three metrics. For example, our MPT models achieve an average gain of 0.98/0.0103/1.27 in BLEU/COMET/BlonDe compared to the CMT-PT models. Our decoding-enhance strategy further enhances the capacity of LLMs, with DeMPT outperforming MPT with an average gain of 0.65/0.0042/0.88. Compared to G-Trans. (+mBART) or MR-Trans (+mBART), DeMPT also demonstrates either superior or comparable performance across all language pairs.
Finally, we observe a similar performance trend among MT models built upon bloomz-7b1-mt. It also indicates that models built upon llama-2-7b outperform those utilizing bloomz-7b1-mt, suggesting that llama-2-7b serves as a more robust foundation model for translation tasks.
| MT-LoRA | MT-PT/CMT-PT | MPT/DeMPT | |
@@ 4 Discussion
In this section, we use bloomz-7b1-mt as the foundation model to discuss our approach.151515Considering page limitation and the consumption of GPUs resources and training time, we use the ZHEN task as a representative to report the BLEU and BlonDe scores. See Appendix DM for further discussions.
@@ 4.1 Effect of Length of Inter-sentence Context
For efficient training, we define the inter-sentence context in Section LABEL:sec:appro as previous sentences with a total tokens not exceeding 256. We are curious about the potential impact of inter-sentence length on the performance of our approach. Consequently, we extend the inter-sentence context length from 256 to 1024 and assess the performance of our approach in the ZHEN task. Figure 4 shows the performance trend of the CMT-PT model and our DeMPT model. As the length of the inter-sentence context increases, both models exhibit a slight enhancement in both BLEU and BlonDe scores. Interestingly, our model with a 256-token inter-sentence context outperforms the CMT-PT model with a 1024-token inter-sentence context in both BLEU and BlonDe scores. This further suggests the effectiveness of our approach in harnessing the capabilities of LLMs for context-aware NMT compared to the concatenation strategy.
4.2 Effect of Multi-phase Strategy
Our multi-phase strategy divides the whole translation into three phases: phase 1 for encoding inter-sentence, phase 2 for encoding intra-sentence, and phase 3 for decoding current sentence. To assess the effect of multi-phase strategy, we compare its performance with two contrasting strategies: merging the first two phases into one (i.e., Merging 1&2) and and merging all phases into a single one (i.e., Merging 1&2&3).161616When merging them all into one, it equals CMT-PT, i.e., the concatenate strategy. Note that in both the two contrasting strategies, we replace the enhanced next word prediction (Eq. 17) in decoding phase with . Table 4 presents the performance of different phrasing strategies. Comparing Merging 1&2 and Merging 1&2&3, it shows that separating the decoding phrase from the encoding marginally improves the performance in both BLEU and BlonDe. Importantly, the comparison of MPT and Merging 1&2 tells that separating the encoding of inter- and intra sentence achieves more gains across all metrics.
| Model | BLEU | COMET | BlonDe |
| MPT | 31.81 | 0.8601 | 50.22 |
| DeMPT | 32.46 | 0.8649 | 50.62 |
[…]
| Merging 1&2 | 31.01 | 0.8503 | 49.91 |
Meanwhile, we conjecture that another benefit of multi-phasing strategy lies in the robustness to the noise contained in document-level context. To test the conjecture, we replace the original inter-sentence context with a random inter-sentence context, meaning we randomly select some sentences from other documents to serve as the inter-sentence context. As shown in Table 5, the performance of both the Merging 1&2&3 and DeMPT models consistently deteriorates when exposed to random context (w/ rnd. CTX). However, the decline is more pronounced for Merging 1&2&3 than for DeMPT (-2.19/0.0102/1.60 vs -0.90/0.0068/0.91). This suggests that DeMPT, owing to its multi-phase strategy, exhibits more robustness in utilizing inter-sentence context in contrast to Merging 1&2&3.
| Model | BLEU | COMET | BlonDe |
| Merging 1&2&3 | 30.82 | 0.8504 | 49.61 |
| w/ rnd. CTX | 28.63 | 0.8402 | 48.01 |
[…]
| w/ rnd. CTX | 31.56 | 0.8581 | 49.71 |
4.3 Human Evaluation
We use the Direct Assessment (DA) method Graham et al. (2017) to manually assess the quality of translations generated by DeMPT and CMT-PT. In this assessment, human evaluators compare the meaning of the MT output with a human-produced reference translation, working within the same language. Specifically, we randomly select 5 documents with a total of 200 groups of sentences from the ZHEN test set. To avoid potential bias in evaluation, we recruit 6 professional translators and ensure each translation from DeMPT or CMT-PT is scored twice by two translators. Table 6 shows the DA scores for CMT-PT and DeMPT. Our DeMPT outperforms CMT-PT by 7.14 DA score, providing strong evidence for the effectiveness of our approach. Further details and results regarding the DA can be found in Appendix H.
| Model | Score_1 | Score_2 | Average |
| CMT-PT | 79.00 | 80.17 | 79.59 |
| DeMPT | 86.17 (+7.17) | 87.30 (+7.13) | 86.73 (+7.14) |
5 Related Work
Due to limited space, we omit the discussion on conventional context-aware MT, focusing instead on LLM-based context-aware MT and prompt tuning for LLMs. Besides, considering our DeMPT’s inspiration from MSP Tan et al. (2022), we offer further discussion on their differences.
LLM-based Context-aware Machine Translation.
While traditional context-aware neural machine translation (NMT) has seen considerable progress in recent years Jean et al. (2017); Wang et al. (2017); Voita et al. (2018); Maruf et al. (2019); Kang et al. (2020); Bao et al. (2021); Sun et al. (2022); Bao et al. (2023), the effective integration of large language models (LLMs) to model inter-sentence context and enhance context-aware translation remains an area of limited exploration. Existing studies mainly focus on the assessment of LLMs’ ability in discourse modeling. For example, Wang et al. (2023) approach context-aware NMT as a task involving long sequence generation, employing a concatenation strategy, and conduct comprehensive evaluations of LLMs such as ChatGPT and GPT-4. Their focus includes the impact of context-aware prompts, comparisons with translation models, and an in-depth analysis of discourse modeling ability. Similarly, Karpinska and Iyyer (2023) engage professional translators to evaluate LLMs’ capacity in context-aware NMT. In contrast, Wu et al. (2024) compare the effectiveness of various parameter-efficient fine-tuning methods on moderately-sized LLMs for context-aware NMT. Besides, Wu and Hu (2023) explore the prompt engineering with GPT language models specifically for document-level (context-aware) MT while Li et al. (2024) experiment with combining sentence-level and document-level translation instructions of varying lengths to fine-tune LLMs. Differently, koneru_etal_2024_contextual propose a post-editing approach to enhance LLMs’ capacity in utilization of inter-sentence context in document-level MT.
Prompt Tuning for Large Language Model.
Liu et al. (2021) and Li and Liang (2021) propose to make LLMs adapt to various tasks by adding trainable prompts (also called continuous prompts) to the original input sequences. In this paradigm, only the continuous prompts are updated during training. Liu et al. (2022) further introduces deep prompt tuning, extending the idea by inserting trainable prompts into all layers of LLMs, rather than just the embedding layer. While these approaches provide a general framework, we focus on enhancing LLM performance specifically for inter-sentence context modeling in context-aware NMT.
Discussion with MSP.
Tan et al. (2022) propose a multi-phase tuning approach (MSP) to enhance the sentence-level translation performance of a multilingual GPT. Our DeMPT mainly differs from MSP in the following aspects: 1) DeMPT adopts a phase-aware prompt to enable distinctive modeling for different inputs, namely inter-sentence contexts, intra-sentence contexts, and the target sentence, a feature not present in MSP; 2) DeMPT incorporates a decoding-enhanced strategy to further improve the effectiveness of utilizing different context information, a capability not available in MSP; 3) DeMPT is designed to alleviate discourse problems in context-aware LLM-based machine translation tasks, rather than addressing sentence-level machine translation tasks as in the case of MSP; 4) DeMPT is designed to adapt LLMs rather than smaller pre-trained model used in MSP.
6 Conclusion
Limitations
Owing to resource limitations, our work is restricted to moderate-scale LLMs, specifically those with 7 billion parameters, and a confined window size of inter-sentence context. It is imperative to acknowledge that the results of our research may differ when employing larger models and extended window sizes for inter-sentence contexts. Considering that English text forms the main body of the training data for LLMs, this paper only focuses on the English-centric translation tasks. The results of non-English-centric translation tasks may vary. We acknowledge these limitations and consider them as avenues for future exploration. Besides, following the finding of Bao et al. (2021), we focus solely on the source-side inter-sentence context in this work. We will explore more about the integration of target-side inter-sentence context in the future.
Appendix A Datasets
Splitting, Preprocessing and Statistics of Datasets.
For all translation tasks, we randomly select 80% document pairs from the corpus as the training set. Both the test set and validation set include 150 document pairs each, randomly sampled from the remaining 20% of document pairs in the corpus. Regarding sentence preprocessing across all datasets for LLM-based models, we segment the sentences with the tokenizer from the respective foundation model. No additional preprocessing steps are performed. For encoder-decoder Transformer models, we segment the source and target sentences into sub-words by a BPE model with 30K merged operations sennrich_etal_acl_2016. We provide the detailed statistic in Table 7. Datasets are downloaded from https://data.statmt.org/news-commentary/v18.
| Dataset | ZHEN | FREN | DEEN | ESEN | RUEN | |||||
| #Doc | #Sent | #Doc | #Sent | #Doc | #Sent | #Doc | #Sent | #Doc | #Sent | |
| Test | 150 | 5,747 | 150 | 5,795 | 150 | 5,967 | 150 | 5,819 | 150 | 5,619 |
Appendix B Training Details
For all encoder-decoder Transformer NMT models, we use the transformer-base setting as in Vaswani et al. (2017), where the learning rate is set to with an inverse-square schedule and warmup steps of , and use Adam optimizer with and . For the other special training settings in G-Trans and MR-Trans, we keep consistent with that provided in their paper. All Transformer NMT models are trained on 4 NVIDIA V100 32GB GPUs with a batch size of 4096. For the models with prompt tuning in Section LABEL:sec:exp, including MT-PT, CMT-PT, MPT and DeMPT models, the length of the trainable prompt is set as 64. During both training and inference, the model generates only the current target sentence, operating in a many-to-one translation mode. For all fine-tuning models in this paper, we set the training epoch to 4, and the warm-up rate to 0.1. We use the log learning rate decay strategy with a maximum learning rate of 5e-5. We collate a mini-batch by counting the total tokens inside the batch and set the batch size as 4096. All fine-tuning models are trained on 4 NVIDIA A800 GPUs with Deespeed Zero 2 offload setting Rajbhandari et al. (2020).171717https://github.com/microsoft/DeepSpeed
Appendix C Traditional Context-aware Models
In this paper, we implement G-Transformer (G-Trans) Bao et al. (2021) and Multi-Resolution Transformer (MR-Trans) Sun et al. (2022) as representatives of traditional context-aware models for comparison. For ease of understanding, we provide a brief introduction to these two models in this section.
G-Transformer.
The transformer model equips a group attention on the lower encoder/decoder layer and a combined attention on the top encoder/decoder layer. For a sentence being translated with its inter-sentence context, the group attention helps maintain locality bias by focusing on intra-sentence context. Meanwhile, the combined attention effectively integrates boundary information, enhancing the translation process with inter-sentence context.
Multi-Resolution Transformer.
The Transformer model does not include any additional modules specifically for modeling inter-sentence context. Instead, it only uses a mixed training set that comprises both sentence-level and document-level instances with varying numbers of sentences. Training on this mixed set allows the Transformer model to handle both sentence-level and document-level translation tasks. In this paper, we implement its Document-to-Sentence variant, which uses all preceding contexts as the source and the current sentence as the target.
Appendix D Comparison of Inference Speed
@@
Appendix E Effect of Prompt Length
As our approach is implemented based on deep prompt tuning, next we compare the impact of the trainable prompt length for MT-PT, CMT-PT, and our DeMPT. Figure 7 shows the performance curves when increasing the prompt length from 32 to 128. We observe that increased prompt length tends to enhance performance for both BLEU and BlonDe, yet the gains exhibit diminishing returns. This finding is consistent with that in Li and Liang (2021); Lester et al. (2021); Tan et al. (2022). We also observe that DeMPT with a prompt length of 64 outperforms both MT-PT and CMT-PT with a prompt length of 128 on both metrics, suggesting the superiority of our approach over the concatenation strategy in enhancing LLMs’ capacity for context-aware NMT.
Appendix F Performance on Contrastive Test Set
We evaluate the models’ ability to resolve discourse inconsistencies using the contrastive test set proposed by Voita et al. (2019a), which focuses on four discourse phenomena such as deixis, lexicon consistency (lex.c), ellipsis inflection (ell.infl), and verb phrase ellipsis (ell.VP) in EnglishRussian translation. Within the test set, each instance comprises a positive translation and several negative ones that vary by only one specific word. The purpose of the contrastive test set is to assess whether a model is more inclined to generate a correct translation as opposed to incorrect variations. Table 9 lists the accuracy of translation prediction on the contrastive test set for MT-PT, CMT-PT and DeMPT. Compared to the context-agnostic MT-PT model, both context-aware CMT-PT and DeMPT models show substantial improvements across the four discourse phenomena. Additionally, DeMPT demonstrates the best performance, surpassing CMT-PT by an average accuracy margin of 3.8.
@@ Appendix F Performance on Contrastive Test Set
Appendix G Effect of Various Contexts for Decoding-enhanced Strategy
| Model | BLEU | COMET | BlonDe |
| MT-PT | 30.99 | 0.8520 | 49.48 |
| CMT-PT | 30.82 | 0.8504 | 49.61 |
[…]
| w/o | 32.33 | 0.8629 | 50.29 |
| w/o | 32.11 | 0.8641 | 50.51 |
We conduct an ablation study on the ZH-EN translation direction using the bloomz-7b-mt model as the foundation model to clarify the effect of the three probabilities in Equation 17, i.e., the effect of various contexts for the heuristic decoding-enhanced strategy. From the Table 10, we observe that removing , i.e., w/o , leads to a significant degradation in the discourse-related metric, namely the BlonDe. This is because the integration enhances the utilization of the inter-sentence context during the decoding phase. We are additionally, removing results in the most substantial degeneration in BLEU metric. This observation demonstrates that our heuristic decoding-enhanced strategy can distinctively improve the utilization of various contexts during the decoding phase.
Appendix H Details of Human Evaluation